OpenStudy (anonymous):

Consider the function, f(x)= {5|x| + 6,if x > 0 \[{5|x| - 6,if x \le 0}\] Evaluate f(-x), if x=10

OpenStudy (anonymous):

Theres no function

OpenStudy (anonymous):

{5|x| + 6,if x > 0 \[{5|x| - 6,if x \le 0}\] Evaluate f(-x), if x=10

OpenStudy (anonymous):

@xXMagicWalrusXx Just submitted teh function...

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

f(-(10))=f(-10)= 5|-10| - 6 = 50-6 = 44

OpenStudy (anonymous):

@dpaInc x=10 right, that means x>0 so won't we use f(x)= {5|x| + 6,if x > 0 this piece of function?

OpenStudy (anonymous):

44

OpenStudy (anonymous):

@himanshu31 same question, we put x=10 right, that means x>0 so won't we use f(x)= {5|x| + 6,if x > 0 this piece of function?

OpenStudy (anonymous):

if we put x=10 we have to find f(-10)..so use x<=10 function

OpenStudy (anonymous):

x<=0

OpenStudy (anonymous):

That means we have to use that piece of function, which deems the value of x after putting in f(x) ??

ganeshie8 (ganeshie8):

nice catch @himanshu31 :)

OpenStudy (anonymous):

Is it?

OpenStudy (anonymous):

@Zeerak yes.. lets say u have to evaluate f(2x-9) for x=10 in this case it becomes f(2*10-9) i.e f(11) ..so u have to use x>0 function

ganeshie8 (ganeshie8):

hey sorry folks.. whatever @dpaInc did is correct.... just went through his reply :P

OpenStudy (anonymous):

dont go what is sign of x ..check what it becomes afters ubstituting

OpenStudy (anonymous):

Okay, I got it now... Thank you so very much Himanshu for the help! :)

OpenStudy (anonymous):

here x=10 so x>0, the function then becomes 5x+6 we need to knw what is f(-x) so it beomes -5x+6.put now x=10 and its done all it matters here is which function is considered is it 5x+6 or the other one